OpenStudy (anonymous):

find the equation of the tangent line to the curve y = 4throot(x) at the point (1,1)

OpenStudy (anonymous):

 At point (1,1) , slope ,m= 1/4 Now equation: y-y1= m (x-x1) y-1 = 1/4(x-1) Just simplify it now.